OpenStudy (anonymous):

(cos6x+6cos4x+15cos2x+10)/cos5x+5cos3x+10cosx

OpenStudy (anonymous):

Do you mean \[\frac{\cos^6x+6\cos^4x+15\cos^2x+10}{\cos^5x+5\cos^3x+10\cos x}~~?\] or \[\frac{\cos6x+6\cos4x+15\cos2x+10}{\cos5x+5\cos3x+10\cos x}~~?\] or \[\frac{\cos^6x+6\cos^4x+15\cos^2x+10}{\cos^5x}+5\cos^3x+10\cos x~~?\] or \[\frac{\cos6x+6\cos4x+15\cos2x+10}{\cos5x}+5\cos3x+10\cos x~~?\]